OpenStudy (anonymous):

solve the equation 3^4x=9^x-10

OpenStudy (anonymous):

\[3^{4x} = 9^{x-10}\\ ( 4x)\ln3=(x-10 )\ln 9\\ 4xln3 = xln9 - 10\ln9\\ \\\ 10 \ln9 = x \ln9 -4xln3 \\ x = \frac{ 10 \ln9 }{ \ln9- 4\ln3 } \\\\\ x \approx-10\]